. Solve π‘₯ 2 𝑑 2𝑦 𝑑2π‘₯ βˆ’ 2 𝑑𝑦 𝑑π‘₯ βˆ’ 4𝑦 = 0Β 


1
Expert's answer
2021-01-31T16:04:47-0500

"\displaystyle
x^2\frac{\mathrm{d}^2y}{\mathrm{d}x^2} - 2x\frac{\mathrm{d}y}{\mathrm{d}x} - 4y = 0\\

\textsf{Let}\,\, y = f(\ln{x})\\

\frac{\mathrm{d}y}{\mathrm{d}x} = \frac{f'(\ln{x})}{x}\\

\begin{aligned}
\frac{\mathrm{d}^2y}{\mathrm{d}x^2} &= -\frac{f'(\ln{x})}{x^2} + \frac{f''(\ln{x})}{x^2}
&= \frac{1}{x^2}\left(f''(\ln{x} - f'(\ln{x})\right)
\end{aligned} \\


x^2\left(\frac{1}{x^2}\left(f''(\ln{x}) - f'(\ln{x})\right)\right) - 2x\cdot\frac{f'(\ln{x})}{x} - 4f(\ln{x}) = 0\\


f''(\ln{x}) - f'(\ln{x} - 2f'(\ln{x}) - 4f(\ln{x}) = 0\\

f''(\ln{x}) - 3f'(\ln{x}) - 4f(\ln{x}) = 0\\


\textsf{Let}\,\, t(l) = f(\ln{x})\\

t'' - 3t' - 4t = 0\\
 
\textsf{Auxiliary equation}; \,\,m^2 - 3m - 4 = 0\\

m = -1, 4\\

t = Ae^{-l} + Be^{4l} \\

t = Ae^{-\ln{x}} + Be^{4\ln{x}} \\

t(l) = f(\ln{x}) = y = \frac{A}{x} + Bx^4 \\

\therefore y = \frac{A}{x} + Bx^4"